# Stochastic Calculus for Pathwise Observables of Markov-Jump Processes: Unification of Diffusion and Jump Dynamics

### v1

@@ III.2 Pathwise Observables: Currents and Densities
Similarly, a density assigns each state $i$ a weight $V_{i}(\tau)$ and then integrates these along the trajectory. These weights can be combined into the state function vector $\bm{V}(\tau)=(V_{1}(\tau),V_{2}(\tau),\dots)^{T}$. With $\bm{V}(\tau)$ we define the density [61]
$\rho_{t}=\int_{\tau=0}^{\tau=t}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau),$ (49)
where $\mathrm{d}\bm{\tau}(\tau)=(\mathrm{d}\tau_{1}(\tau),\mathrm{d}\tau_{2}(\tau),\dots)^{T}$.
@@
Figure 3: Example trajectory and examples of an arising current and density. The part of the trajectory in (a) which contributes to the density $\rho_{\tau}$ with state function, $\bm{V}^{T}=(0,1,0)$, is marked in blue. Similarly, the transitions contributing to the current $J_{\tau}$ with transition weights, $\kappa_{ij}=\delta_{i2}\delta_{j3}-\delta_{i3}\delta_{j2}$, are marked by orange arrows with numbers representing the weight of the transition. The resulting density and current can be seen in (b) and (c), respectively.
@@ III.3 From Stratonovich Increments to Observable Correlations
$\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}U_{xy}^{\alpha}\sum_{i,j}U^{\beta}_{ij}[\cdot],$ (50b)
where $\cdot$] in the first line denotes functions which depend on $\bm{z}$, $\bm{z}^{\prime}$, $\tau$, and $\tau^{\prime}$. Similarly, the functions entering $\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ depend on $x$, $y$, $i$, $j$, $\tau$, and $\tau^{\prime}$. The functions $M(\bm{z})$ and $N(\bm{z})$ may be scalar- or vector-valued functions, while $U^{\alpha}$ and $U^{\beta}$ are two-point quantities, e.g., $\bm{\kappa}$. Equation (50b) corresponds to Eq. (50a) where each continuous spatial integral has been changed to a discrete double sum. While one may expect the integrals to be substituted by a single sum, the double sum is indeed required to include non-local quantities, e.g., currents.
We will now consider two observables $X_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c),k}},\rho_{t}^{\mathrm{(c),k}}\}$ and $Y_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c),l}},\rho_{t}^{\mathrm{(c),l}}\}$, where the superscript is used to distinguish discrete from the continuous setting and $k,l\in\{1,2\}$ is the index for currents $J_{t}^{\mathrm{(c),1}},J_{t}^{\mathrm{(c),2}}$ and densities $\rho_{t}^{\mathrm{(c),1}},\rho_{t}^{\mathrm{(c),2}}$. These currents and densities may have different current-/density-defining “window” functions, which we take into consideration here.
@@
where $m=1$ for $X_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},l}$, w.l.o.g. $m=2$ for $X_{t}^{\mathrm{(c)}}=J_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},l}$, and $m=3$ for $X_{t}^{\mathrm{(c)}}=J_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=J_{t}^{\mathrm{(c)},l}$. For density-density covariance ($m=1$), the increment correlations read [96]
${\Xi}_{1}^{\bm{z}\bm{z}^{\prime}}$ $=\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}+\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}},$ (52)
@@
$\mathrm{cov}({\rho}_{t}^{\mathrm{c},k},{\rho}_{t}^{\mathrm{c},l})=$ $\hat{\mathcal{I}}^{t,\mathrm{c}}_{V^{k},V^{l}}\left[P_{\mathrm{s}}(\bm{z},\tau;\bm{z}^{\prime},\tau^{\prime})+P_{\mathrm{s}}(\bm{z}^{\prime},\tau;\bm{z},\tau^{\prime})-p_{\mathrm{s}}(\bm{z})p_{\mathrm{s}}(\bm{z}^{\prime})\right],$
$\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})+P_{\mathrm{s}}(x,\tau^{\prime};i,\tau)-p_{x}^{\mathrm{s}}p_{i}^{\mathrm{s}}\right].$ (53)
@@
which include correlations of a Stratonovich increment $\circ\mathrm{d}\bm{x}_{\tau}$ and $\mathrm{d}n_{xy}(\tau)$ with a future event. The evaluation of such correlations is possible by means of the generalized time-reversal symmetry, i.e., dual-reversal symmetry [141, 133, 96, 125]. In both cases, the dual system (denoted by ) has the same steady state probability whereas the irreversible drift $\bm{j}_{\mathrm{s}}(\bm{x})$ has the opposite sign $\bm{j}^{\ddagger}_{\mathrm{s}}(\bm{x})=-\bm{j}_{\mathrm{s}}(\bm{x})$. It is useful to define current and dual current operators; in the continuous-space setting, these are [133, 96, 125] $\hat{\bm{j}}_{\bm{x}}=p_{s}(\bm{x})^{-1}\bm{j}_{\mathrm{s}}(\bm{x})-p_{\mathrm{s}}(\bm{x})\bm{D}\nabla_{\bm{x}}p_{\mathrm{s}}(\bm{x})^{-1}$ and $\hat{\bm{j}}_{\bm{x}}^{\ddagger}=p_{s}(\bm{x})^{-1}\bm{j}_{\mathrm{s}}(\bm{x})+p_{\mathrm{s}}(\bm{x})\bm{D}\nabla_{\bm{x}}p_{\mathrm{s}}(\bm{x})^{-1}$. The corresponding operators for discrete-state systems are much simpler and read $\hat{j}_{xy}=r_{xy}$ and $\hat{j}_{xy}^{\ddagger}=r_{yx}p_{y}^{\mathrm{s}}/p_{x}^{\mathrm{s}}$, respectively. The latter is the dual current operator reversing the steady-state currents [6] . When these current operators act on $p_{x}^{\mathrm{s}}$, they result in the probability flux and dual probability flux on edge $x\to y$, i.e., $\hat{j}_{xy}p^{\mathrm{s}}_{x}=r_{xy}p_{x}^{\mathrm{s}}=J_{xy}^{\mathrm{s}}$ and $\hat{j}^{\ddagger}_{xy}p^{\mathrm{s}}_{x}=r_{yx}p_{y}^{\mathrm{s}}=J_{yx}^{\mathrm{s}}$. By means of the flux operators, it can be shown that
@@
$\mathrm{cov}({J}_{t}^{\mathrm{c},k},{\rho}_{t}^{\mathrm{c},l})$ $=\hat{\mathcal{I}}^{t,\mathrm{c}}_{\bm{U}^{k},V^{l}}\left[\hat{\bm{j}}_{\bm{z}^{\prime}}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})+\hat{\bm{j}}^{\ddagger}_{\bm{z}}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})-\bm{j}_{\mathrm{s}}(\bm{z})p_{\mathrm{s}}(\bm{z}^{\prime})\right],$
$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})$ $=\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\hat{j}_{xy}P(x,t^{\prime}|i)p_{i}^{\mathrm{s}}+\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}p_{i}^{\mathrm{s}}\right].$ (58)
@@
$\mathrm{cov}({J}_{t}^{\mathrm{c},k},J_{t}^{\mathrm{c},l})=$ $2t\int\mathrm{d}\bm{z}(\bm{U}^{k}(\bm{z}))^{T}\bm{D}\bm{U}^{l}(\bm{z})p_{\mathrm{s}}(\bm{z})$
@@
$\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $t\sum_{x,y}\kappa_{xy}^{k}\kappa_{xy}^{l}r_{xy}p_{x}^{\mathrm{s}}+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\hat{j}_{xy}\hat{j}_{ji}^{\ddagger}P(x,t^{\prime}|j)p_{j}^{\mathrm{s}}+\hat{j}_{ij}\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}J^{\mathrm{s}}_{ij}\right],$ (60)
@@ III.4 Transient Covariances
$\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}U_{xy}^{\alpha}(\tau)\sum_{i,j}U^{\beta}_{ij}(\tau^{\prime})[\cdot],$ (61)
@@
$\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P(x,\tau;i,\tau^{\prime})-p_{x}(\tau)p_{i}(\tau^{\prime})\right],$
$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime})+\mathds{1}_{\tau<\tau^{\prime}}r_{yx}P(i,\tau^{\prime}|y,\tau)p_{y}(\tau)-r_{xy}p_{x}(\tau)p_{i}(\tau^{\prime})\right],$
$\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}\kappa_{xy}^{k}(\tau)\kappa_{xy}^{l}(\tau)r_{xy}p_{x}(\tau)$
$+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}r_{ij}P(x,\tau|j,\tau^{\prime})p_{i}(\tau^{\prime})+\mathds{1}_{\tau<\tau^{\prime}}r_{ij}r_{xy}P(i,\tau^{\prime}|y,\tau)p_{y}(\tau)-r_{xy}p_{x}(\tau)r_{ij}p_{i}(\tau^{\prime})\right],$ (62)

### v2

@@ III.2 Pathwise Observables: Currents and Densities
Similarly, a density assigns each state $i$ a weight $V_{i}(\tau)$ and then integrates these along the trajectory. These weights can be combined into the state function $V_{\tau}\equiv V(x_{\tau},\tau)=\sum_{k}\delta_{x_{\tau}k}V_{k}(\tau)$. With ${V}_{\tau}$ we define the density [38]
[…]
@@
[…]
@@ III.3 From Stratonovich Increments to Observable Correlations
$\hat{{I}}^{t}_{{\color[rgb]{0,0,0}G}^{\alpha},{\color[rgb]{0,0,0}G}^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}{\color[rgb]{0,0,0}G}_{xy}^{\alpha}\sum_{i,j}{\color[rgb]{0,0,0}G}^{\beta}_{ij}[\cdot],$ (58b)
where $\cdot$] in the first line denotes functions which depend on $\bm{z}$, $\bm{z}^{\prime}$, $\tau$, and $\tau^{\prime}$. Similarly, the functions entering $\hat{{I}}^{t}_{{\color[rgb]{0,0,0}G}^{\alpha},{\color[rgb]{0,0,0}G}^{\beta}}[\cdot]$ depend on $x$, $y$, $i$, $j$, $\tau$, and $\tau^{\prime}$. The functions $M(\bm{z})$ and $N(\bm{z})$ may be scalar- or vector-valued functions, while ${\color[rgb]{0,0,0}G}^{\alpha}$ and ${\color[rgb]{0,0,0}G}^{\beta}$ are two-point quantities, e.g., $\bm{\kappa}$. Equation (58b) corresponds to Eq. (58a) where each continuous spatial integral has been changed to a discrete double sum. While one may expect the integrals to be substituted by a single sum, the double sum is indeed required to include non-local quantities, e.g., currents.
We will now consider two observables $X_{t}^{\mathrm{(c)}}\in\{J_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}},\rho_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}\}$ and $Y_{t}^{\mathrm{(c)}}\in\{J_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}},\rho_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}\}$, where the superscript (c) is used to distinguish discrete (no superscript) from the continuous (superscript $\mathrm{c}$) setting and $k,l\in\{1,2\}$ is the index for currents $J_{t,{\color[rgb]{0,0,0}1}}^{\mathrm{(c)}},J_{t,{\color[rgb]{0,0,0}2}}^{\mathrm{(c)}}$ and densities $\rho_{t,{\color[rgb]{0,0,0}1}}^{\mathrm{(c)}},\rho_{t,{\color[rgb]{0,0,0}2}}^{\mathrm{(c)}}$. These currents and densities may have different current-/density-defining functions, which we take into consideration here.
@@
where the $\Xi$ matrices are defined through increment correlations, see Eqs. (60), (62), and (67). The index $m$ we use to distinguish between different observable correlations: $m=1$ for $X_{t}^{\mathrm{(c)}}=\rho_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}$, w.l.o.g. $m=2$ for $X_{t}^{\mathrm{(c)}}=J_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}$, and $m=3$ for $X_{t}^{\mathrm{(c)}}=J_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}$ and $Y_{t}^{\mathrm{(c)}}=J_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}$. For density-density covariance ($m=1$), the increment correlations read [36]
${\Xi}_{1}^{\bm{z}\bm{z}^{\prime}}$ $=\langle{\color[rgb]{0,0,0}\mathbb{1}}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}+\langle{\color[rgb]{0,0,0}\mathbb{1}}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}},$ (60)
@@
$\mathrm{cov}({\rho}_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{c}},{\rho}_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{c}})=$ $\hat{\mathcal{I}}^{t,\mathrm{c}}_{V^{k},V^{l}}\left[P_{\mathrm{s}}(\bm{z},\tau;\bm{z}^{\prime},\tau^{\prime})+P_{\mathrm{s}}(\bm{z}^{\prime},\tau;\bm{z},\tau^{\prime})-p_{\mathrm{s}}(\bm{z})p_{\mathrm{s}}(\bm{z}^{\prime})\right],$
$\mathrm{cov}({\rho}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})+P_{\mathrm{s}}(x,\tau^{\prime};i,\tau)-p_{x}^{\mathrm{s}}p_{i}^{\mathrm{s}}\right].$ (61)
@@
which include correlations of a Stratonovich increment $\circ\mathrm{d}\bm{x}_{\tau}$ and $\mathrm{d}n_{xy}(\tau)$ with a future event. The evaluation of such correlations is possible by means of the generalized time-reversal symmetry, i.e., dual-reversal symmetry [73, 37, 36, 35]. In both cases, the dual system (denoted by ) has the same steady state probability whereas the irreversible drift $\bm{j}_{\mathrm{s}}(\bm{x})$ has the opposite sign $\bm{j}^{\ddagger}_{\mathrm{s}}(\bm{x})=-\bm{j}_{\mathrm{s}}(\bm{x})$. It is useful to define current and dual current operators; in the continuous-space setting, these are [37, 36, 35] $\hat{\bm{j}}_{\bm{x}}=p_{s}(\bm{x})^{-1}\bm{j}_{\mathrm{s}}(\bm{x})-p_{\mathrm{s}}(\bm{x})\bm{D}\nabla_{\bm{x}}{\color[rgb]{0,0,0}\left[{\color[rgb]{0,0,0}p_{\mathrm{s}}(\bm{x})^{-1}}\right]}$ and $\hat{\bm{j}}_{\bm{x}}^{\ddagger}=p_{s}(\bm{x})^{-1}\bm{j}_{\mathrm{s}}(\bm{x})+p_{\mathrm{s}}(\bm{x})\bm{D}\nabla_{\bm{x}}{\color[rgb]{0,0,0}\left[{\color[rgb]{0,0,0}p_{\mathrm{s}}(\bm{x})^{-1}}\right]}$. The corresponding operators for discrete-state systems are much simpler and read $\hat{j}_{xy}=r_{xy}$ and $\hat{j}_{xy}^{\ddagger}=r_{yx}p_{y}^{\mathrm{s}}/p_{x}^{\mathrm{s}}$, respectively. The latter is the dual current operator reversing the steady-state currents [144] . When these current operators act on $p_{x}^{\mathrm{s}}$, they result in the probability flux and dual probability flux on edge $x\to y$, i.e., $\hat{j}_{xy}p^{\mathrm{s}}_{x}=r_{xy}p_{x}^{\mathrm{s}}=J_{xy}^{\mathrm{s}}$ and $\hat{j}^{\ddagger}_{xy}p^{\mathrm{s}}_{x}=r_{yx}p_{y}^{\mathrm{s}}=J_{yx}^{\mathrm{s}}$. By means of the flux operators, it can be shown that
@@
$\mathrm{cov}({J}_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{c}},{\rho}_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{c}})$ $=\hat{\mathcal{I}}^{t,\mathrm{c}}_{\bm{U}^{k},V^{l}}\left[\hat{\bm{j}}_{\bm{z}^{\prime}}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})+\hat{\bm{j}}^{\ddagger}_{\bm{z}}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})-\bm{j}_{\mathrm{s}}(\bm{z})p_{\mathrm{s}}(\bm{z}^{\prime})\right],$
$\mathrm{cov}({J}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})$ $=\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\hat{j}_{xy}P(x,t^{\prime}|i)p_{i}^{\mathrm{s}}+\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}p_{i}^{\mathrm{s}}\right].$ (66)
@@
$\mathrm{cov}({J}_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{c}},J_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{c}})=$ $2t\int\mathrm{d}\bm{z}(\bm{U}^{k}(\bm{z}))^{T}\bm{D}\bm{U}^{l}(\bm{z})p_{\mathrm{s}}(\bm{z})$
@@
$\mathrm{cov}(J_{t,{\color[rgb]{0,0,0}k}},{J}_{t,{\color[rgb]{0,0,0}l}})=$ $t\sum_{x,y}\kappa_{xy}^{k}\kappa_{xy}^{l}r_{xy}p_{x}^{\mathrm{s}}+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\hat{j}_{xy}\hat{j}_{ji}^{\ddagger}P(x,t^{\prime}|j)p_{j}^{\mathrm{s}}+\hat{j}_{ij}\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}J^{\mathrm{s}}_{ij}\right],$ (68)
@@ III.4 Transient Covariances
$\hat{{I}}^{t}_{{\color[rgb]{0,0,0}G}^{\alpha},{\color[rgb]{0,0,0}G}^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}{\color[rgb]{0,0,0}G}_{xy}^{\alpha}(\tau)\sum_{i,j}{\color[rgb]{0,0,0}G}^{\beta}_{ij}(\tau^{\prime})[\cdot],$ (69)
@@
$\mathrm{cov}({\rho}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P(x,\tau;i,\tau^{\prime})-p_{x}(\tau)p_{i}(\tau^{\prime})\right],$
$\mathrm{cov}({J}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[{\color[rgb]{0,0,0}\mathbb{1}}_{\tau>\tau^{\prime}}r_{xy}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime})+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}r_{yx}P(i,\tau^{\prime}|y,\tau)p_{y}(\tau)-r_{xy}p_{x}(\tau)p_{i}(\tau^{\prime})\right],$
$\mathrm{cov}(J_{t,{\color[rgb]{0,0,0}k}},{J}_{t,{\color[rgb]{0,0,0}l}})=$ $\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}\kappa_{xy}^{k}(\tau)\kappa_{xy}^{l}(\tau)r_{xy}p_{x}(\tau)$
$+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[{\color[rgb]{0,0,0}\mathbb{1}}_{\tau>\tau^{\prime}}r_{xy}r_{ij}P(x,\tau|j,\tau^{\prime})p_{i}(\tau^{\prime})+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}r_{ij}r_{xy}P(i,\tau^{\prime}|y,\tau)p_{y}(\tau)-r_{xy}p_{x}(\tau)r_{ij}p_{i}(\tau^{\prime})\right],$ (70)